OpenStudy (anonymous):

Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step? A. 2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4 B. 2x + (4 - x) = 8 C. 2x + y = x + y D. 3x + 2y = 12 E. x = 4

OpenStudy (anonymous):

x = 4, y = 0

OpenStudy (anonymous):

So E ?

OpenStudy (anonymous):

nope

OpenStudy (campbell_st):

2x + y = 8 (1) x + y = 4 (2) double equation 2 and subtract it from equation 1 2x + y = 8 - 2x +2y = 8 --------------- -y = 0 so y = 0 sub into equation 2 to find x x = 4

OpenStudy (anonymous):

2x + y = 8--------> 2x + y = 8

OpenStudy (anonymous):

so answer is A

OpenStudy (anonymous):

Alexis, keep asking. I wanna answer XD

OpenStudy (anonymous):

yes, answer is A. this is a Didactic question, just to make sure u understand the steps in eliminating one variable. bear in mind you have to subtract (x-x, y-y) equations in order to obtain a third one with a single variable in it